# VScan: Rethinking Visual Token Reduction for Efficient Large Vision-Language Models

### v2

@@ 5.1 Experimental Settings
Models and Baselines. We applied VScan to four popular LVLMs with different architectures to evaluate its general effectiveness. Specifically, we follow previous work in this field to compare performance on LLaVA-1.5-7B [39], LLaVA-NeXT-7B [40], and Video-LLaVA-7B [37]. Furthermore, we also present experimental results on the recent Qwen-2.5-VL [6] model with various LLM sizes (3B, 7B, 32B). We compare the performance of our approach with 6 state-of-the-art visual token reduction methods: ToMe [7], FastV [12], SparseVLM [69], HiRED [4], PyramidDrop [60], and VisionZip [62].
@@
Benchmarks. We conduct extensive experiments on 9 standard image understanding benchmarks, including visual question answering benchmarks such as GQA [25], ScienceQA [44], VQAv2 [23], TextVQA [52] and VizWiz [24]; and multi-modal reasoning benchmarks such as MMBench [43], MMBench-CN [43], MME [20], and POPE [35]. We also include evaluations on 3 more challenging referring grounding tasks using RefCOCO [30], RefCOCO+ [30], and RefCOCOg [47]. Additionally, we evaluate our approach on 4 video question answering benchmarks: TGIF [27], MSVD [10], MSRVTT [61], and ActivityNet [64].
@@
Implementation Details. We adhere to the default inference settings for each evaluated LVLM as specified in their respective codebases. Additionally, we perform local scan at a shallow layer, specifically at $l=6$ for LLaVA-series models and $l=8$ for Qwen-2.5-VL. For LLM-stage pruning, we select the middle layer as $k=16$ for LLaVA-series models and $k=14$ for Qwen-2.5-VL. By default, we fix the retention rate at the LLM middle layer to $R_{2}=33.3\%$, and adjust $R_{1}$ accordingly to achieve the target average reduction rate. For instance, to achieve an average retention rate of 11.1%, we set $R_{1}=16.7\%$ and $R_{2}=33.3\%$. Note that these design choices are grounded in our analysis in Section 3, and we also provide comprehensive ablation studies in Section 5.3.
@@ 5.2 Results and Discussions
Table 3: Performance comparisons on LLaVA-NeXT-7B [40] across 9 image understanding benchmarks. The best results in each setting are bolded, and the second-best are underlined.
@@
LLaVA-NeXT-7B [40] CVPR’24 64.2 67.4 60.6 1851 86.5 70.1 81.8 61.3 57.6 100.0%
@@
FastV [12] ECCV’24 55.9 61.6 51.9 1661 71.7 62.8 72.9 55.7 53.1 88.7%
HiRED [4] AAAI’25 59.3 64.2 55.9 1690 83.3 66.7 75.7 58.8 54.2 93.9%
PyramidDrop [60] CVPR’25 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 91.4%
VisionZip [62] CVPR’25 59.3 63.1 - 1702 82.1 67.3 76.2 58.9 - 94.0%
@@
Figure 5: Performance comparisons on Qwen-2.5-VL [6] with different LLM sizes (3B/7B/32B) across 3 image understanding benchmarks. We present the performance of different approaches at 4 various retention rates, along with the original model performance without token reduction.
@@
Results on LLaVA-1.5. In Table 2, we apply our approach to LLaVA-1.5-7B [39] and compare its performance with other baselines across 9 image understanding tasks. With only 128 and 192 tokens per image instead of the original 576, our approach nearly retains the performance of the original LLaVA-1.5, with only negligible performance declines of 1.0% and 1.2%, respectively. Our approach becomes even more advantageous with higher reduction rates: With an aggressive 88.9% reduction rate, our approach results in only a 3.3% degradation in average performance across benchmarks, outperforming the second-best VisionZip [62] by a substantial margin of 4.0%. These results suggest that our method effectively maintains high performance while reducing the number of visual tokens processed.
@@
Results on LLaVA-NeXT. In Table 3, we further compare the performance achieved by our approach with other state-of-the-arts on the more advanced LLaVA-NeXT-7B [40]. We present the performance of all approaches under a fixed budget of 320 tokens per image, corresponding to an 88.9% reduction rate. Our approach continues to achieve superior performance on LLaVA-NeXT-7B [40], attaining the best performance on 6 out of 9 benchmarks, and achieving 95.4% of the original LLaVA-NeXT-7B [40] performance with only 11.1% of the token budget without additional training.
@@
Results on Qwen-2.5-VL. To further validate the general effectiveness of our approach, we apply it to the recent Qwen-2.5-VL [6] with three different LLM scales and visualize the performance on three image understanding benchmarks in Figure 5. As shown, our approach consistently outperforms FastV [12] and PDrop [60] across all retention rates (from 11.1% to 50%) and model scales. We also observe that on MME, both FastV and PDrop exhibit degraded performance under low token budgets as the LLM scales from 7B to 32B. We speculate that this is due to larger LLMs introducing stronger language priors, which can hinder the accurate selection of critical visual tokens. In contrast, our two-stage reduction strategy mitigates this bias and remains robust across LLM scales.
@@
Table 4: Performance comparisons on Qwen-2.5-VL-7B [6] across 3 referring expression comprehension benchmarks: RefCOCO, RefCOCO+, and RefCOCOg. The best results in each setting are bolded, and the second-best are underlined. Evaluation is based on our re-implementation.
Method Venue RefCOCO RefCOCO+ RefCOCOg Average
@@
Upper Bound, 4$\sim$16384 Tokens (100%)
Qwen-2.5-VL-7B [6] arXiv’25 89.45 92.56 85.16 83.50 89.02 79.15 86.76 87.24 100.0%
@@
FastV [12] ECCV’24 85.27 87.84 82.28 79.02 82.95 72.86 82.95 83.32 94.8%
PyramidDrop [60] CVPR’25 87.79 91.00 83.22 81.48 86.55 74.02 84.62 85.10 97.2%
@@
FastV [12] ECCV’24 73.85 73.38 74.21 66.75 68.88 62.65 71.06 71.86 81.2%
PyramidDrop [60] CVPR’25 77.52 80.82 72.07 70.27 75.48 63.33 74.86 75.65 85.1%
@@
FastV [12] ECCV’24 43.57 46.81 40.86 39.47 43.78 36.02 43.04 42.69 48.5%
PyramidDrop [60] CVPR’25 46.46 53.83 37.23 42.29 47.76 32.81 45.32 44.91 50.4%
@@
We extend our comparisons to more challenging grounding tasks and present the performance of different approaches in Table 4. Compared to image understanding tasks, these grounding tasks require higher token budgets to preserve visual information necessary for precise localization. A 75% reduction rate, for instance, halves the performance of FastV [12] and PDrop [60]. In this challenging scenario, our approach still robustly maintains 80.7% of the original performance.
Results on Video-LLaVA. In Appendix A.1, we further assess the effectiveness of our approach on video understanding tasks and compare its performance against other approaches on Video-LLaVA-7B [37]. The results illustrate that with a 25% token budget, our approach maintains nearly 100% of the original Video-LLaVA-7B’s performance, consistently outperforming other methods.
@@ 5.3 Ablation Studies
Varying Retention Rates $R_{1}$ and $R_{2}$. We analyze how different retention rate configurations affect performance by varying the retention rates $R_{1}$ and $R_{2}$ during the visual encoding and LLM decoding stages, respectively. As shown in Table 5 (a), we observe that relying solely on token selection in the visual encoder or applying overly aggressive pruning during LLM decoding leads to suboptimal performance. Instead, a more balanced and gradual two-stage pruning strategy, i.e., $R_{1}=16.7\%$ and $R_{2}=33.3\%$, achieves the best performance. These results also validate the effectiveness of combining both visual token reduction strategies to jointly improve efficiency and maintain accuracy.
@@
Different Encoding Layers for Local Scan. In Table 5 (c), we explore the effect of performing local token selection at different layers of the visual encoder. Consistent with our empirical findings in Section 3, we find that applying the local scan at a shallow layer ($l=6$) yields the best performance. However, performing the local scan at very early ($l=2$) or the output layer ($l=23$) leads to a noticeable performance drop, with performance on GQA falling to 57.1 and 57.3, respectively.

### v3

@@ 5.1 Experimental Settings
Models. We evaluate the general effectiveness of VScan by applying it to four popular LVLMs with diverse architectures. Following prior work in this field, we first compare performance on LLaVA-1.5-7B Liu et al. [2024a], a widely adopted academic baseline that maps each image input to 576 tokens, and LLaVA-NeXT-7B Liu et al. [2024b], which improves high-resolution understanding by encoding an image into up to 2,880 visual tokens. We further assess Video-LLaVA-7B Lin et al. [2024], which extends the LLaVA framework to videos, processing up to 8 frames with 2,048 visual tokens. Finally, we are among the first to report experimental results on the recent Qwen-2.5-VL Bai et al. [2025], tested across different LLM sizes (3B, 7B, 32B). This model incorporates dynamic resolution processing to handle images of varying sizes, supporting visual token counts ranging from 4 to 16,384.
@@
Benchmarks and Metrics. We conduct extensive experiments on 9 standard image understanding benchmarks, including visual question answering benchmarks such as GQA Hudson and Manning [2019], ScienceQA Lu et al. [2022], VQAv2 Goyal et al. [2017], TextVQA Singh et al. [2019] and VizWiz Gurari et al. [2018]; multi-modal reasoning benchmarks such as MMBench Liu et al. [2024d], MMBench-CN Liu et al. [2024d], MME Fu et al. [2023], and POPE Li et al. [2023b]. We also include evaluations on 3 more challenging referring grounding tasks using RefCOCO Kazemzadeh et al. [2014], RefCOCO+ Kazemzadeh et al. [2014], and RefCOCOg Mao et al. [2016], and report the accuracy achieved by different approaches. In these grounding tasks, a localization is considered correct if the predicted bounding box has an IoU score of at least 0.5 with the ground truth. Additionally, we evaluate our approach on 4 video question answering benchmarks: TGIF Jang et al. [2017], MSVD Chen and Dolan [2011], MSRVTT Xu et al. [2016], and ActivityNet Yu et al. [2019]. We follow previous work Maaz et al. [2024]; Xing et al. [2025] to utilize both accuracy and the ChatGPT score [^2] Evaluated using gpt-3.5-turbo: https://platform.openai.com/docs/models/gpt-3.5-turbo. as key performance metrics for these video-based benchmarks. The evaluation protocol and prompts are detailed in Section B.4.
Baselines. We compare the performance of our approach with 6 state-of-the-art visual token pruning methods: ToMe Bolya et al. [2023], FastV Chen et al. [2024a], SparseVLM Zhang et al. [2025c], HiRED Arif et al. [2025], PyramidDrop Xing et al. [2025], and VisionZip Yang et al. [2025]. (1) ToMe Bolya et al. [2023], which uses bipartite soft matching to iteratively merge similar tokens within ViT layers; (2) FastV Chen et al. [2024a], which drops visual tokens in early layers of the LLM, guided by text-oriented attention score; (3) SparseVLM Zhang et al. [2025c], which selects vision-relevant text tokens to evaluate the significance of visual tokens; (4) HiRED Arif et al. [2025], which dynamically assigns token budgets to sub-images for high-resolution image inputs; (5) PyramidDrop Xing et al. [2025], which divides the LLM into stages and drops a portion of visual tokens at the end of each stage; (6) VisionZip Yang et al. [2025], which selects a set of dominant tokens and merges unselected tokens contextually. To ensure a fair comparison, we directly report the results of these baselines from their respective original papers unless stated otherwise.
@@
Implementation Details. We adhere to the default inference settings for each evaluated LVLM as specified in their respective codebases. Additionally, we perform local scan at a shallow layer, specifically at $l=6$ for LLaVA-series models and $l=8$ for Qwen-2.5-VL. For LLM-stage pruning, we select the middle layer as $k=16$ for LLaVA-series models and $k=14$ for Qwen-2.5-VL-7B. For Qwen-2.5-VL-3B and -32B, we similarly select the exact middle layer for pruning. By default, we fix the retention rate at the LLM middle layer to $R_{2}=33.3\%$, and adjust $R_{1}$ accordingly to achieve the target average reduction rate. For instance, to achieve an average retention rate of 11.1%, we set $R_{1}=16.7\%$ and $R_{2}=33.3\%$. Note that these design choices are grounded in our analysis in Section 3, and we also provide comprehensive ablation studies in Section 5.3. Code will be made publicly available upon acceptance to facilitate reproducibility.
@@ 5.2 Results and Discussions
Table 2: Performance comparisons on LLaVA-1.5-7B Liu et al. [2024a] across 9 image understanding benchmarks. The best results in each setting are bolded, and the second-best are underlined.
[…]
Upper Bound, 576 Tokens (100%), 3.817 TFLOPs
LLaVA-1.5-7B Liu et al. [2024a] CVPR’24 61.9 64.7 58.1 1862 85.9 69.5 78.5 58.2 50.0 100.0%
Retain 192 Tokens in Average ($\downarrow$ 66.7%), $\sim$1.253 TFLOPs
ToMe Bolya et al. [2023] ICLR’23 54.3 60.5 - 1563 72.4 65.2 68.0 52.1 - 88.5%
FastV Chen et al. [2024a] ECCV’24 52.7 61.2 57.0 1612 64.8 67.3 67.1 52.5 50.8 90.4%
SparseVLM Zhang et al. [2025c] ICML’25 59.5 64.1 53.7 1787 85.3 68.7 75.6 57.8 50.5 97.6%
PyramidDrop Xing et al. [2025] CVPR’25 57.3 63.3 56.8 1797 82.3 69.0 75.1 56.5 51.1 97.2%
VisionZip Yang et al. [2025] CVPR’25 59.3 63.0 - 1783 85.3 68.9 77.4 57.3 - 97.8%
VScan (Ours) - 60.6 63.9 57.4 1806 86.2 68.6 77.8 57.7 50.4 99.0%
Retain 128 Tokens in Average ($\downarrow$ 77.8%), $\sim$0.833 TFLOPs
ToMe Bolya et al. [2023] ICLR’23 52.4 53.3 - 1343 62.8 59.6 63.0 49.1 - 80.4%
FastV Chen et al. [2024a] ECCV’24 49.6 56.1 56.4 1490 59.6 60.2 61.8 50.6 51.3 85.4%
SparseVLM Zhang et al. [2025c] ICML’25 58.4 64.5 51.1 1746 85.0 68.6 73.8 56.7 51.4 96.4%
PyramidDrop Xing et al. [2025] CVPR’25 57.1 61.6 56.6 1761 82.3 68.4 72.9 56.6 51.0 96.2%
VisionZip Yang et al. [2025] CVPR’25 57.6 62.0 - 1763 83.2 68.9 75.6 56.8 - 96.2%
VScan (Ours) - 59.8 63.0 58.0 1792 86.1 68.9 77.1 57.3 51.7 98.8%
Retain 64 Tokens in Average ($\downarrow$ 88.9%), $\sim$0.415 TFLOPs
ToMe Bolya et al. [2023] ICLR’23 48.6 43.7 - 1138 52.5 50.0 57.1 45.3 - 70.1%
FastV Chen et al. [2024a] ECCV’24 46.1 48.0 52.7 1256 48.0 51.1 55.0 47.8 50.8 76.7%
SparseVLM Zhang et al. [2025c] ICML’25 53.8 60.1 46.1 1589 77.5 69.8 68.2 53.4 50.1 90.4%
PyramidDrop Xing et al. [2025] CVPR’25 47.5 58.8 50.5 1561 55.9 69.2 69.2 50.6 50.7 86.6%
VisionZip Yang et al. [2025] CVPR’25 55.1 60.1 - 1690 77.0 69.0 72.4 55.5 - 92.7%
VScan (Ours) - 58.3 62.1 55.7 1698 85.0 69.1 75.4 55.6 51.8 96.7%
Table 3: Performance comparisons on LLaVA-NeXT-7B Liu et al. [2024b] across 9 image understanding benchmarks. The best results in each setting are bolded, and the second-best are underlined.
@@
LLaVA-NeXT-7B Liu et al. [2024b] CVPR’24 64.2 67.4 60.6 1851 86.5 70.1 81.8 61.3 57.6 100.0%
@@
FastV Chen et al. [2024a] ECCV’24 55.9 61.6 51.9 1661 71.7 62.8 72.9 55.7 53.1 88.7%
HiRED Arif et al. [2025] AAAI’25 59.3 64.2 55.9 1690 83.3 66.7 75.7 58.8 54.2 93.9%
PyramidDrop Xing et al. [2025] CVPR’25 56.4 63.4 56.2 1663 77.6 67.5 73.5 54.4 54.1 91.4%
VisionZip Yang et al. [2025] CVPR’25 59.3 63.1 - 1702 82.1 67.3 76.2 58.9 - 94.0%
@@
Figure 5: Performance comparisons on Qwen-2.5-VL Bai et al. [2025] with different LLM sizes (3B/7B/32B) across 3 image understanding benchmarks. We present the performance of different approaches at 4 various retention rates, along with the original model performance without token reduction.
@@
Results on LLaVA-1.5. In Table 2, we apply our approach to LLaVA-1.5-7B Liu et al. [2024a] and compare its performance with other baselines across 9 image understanding tasks. Following previous work Zhang et al. [2025c]; Yang et al. [2025] in this field, we present a comparative analysis of performance across three settings, retaining an average of 192, 128, and 64 visual tokens across all 32 layers in LLaMA Touvron et al. [2023], corresponding to reduction rates of 66.7%, 77.8%, and 88.9%, respectively. We also calculate and present the percentage of performance retention in the final column, using the vanilla LLaVA-1.5 performance as the 100% upper limit. As shown in the table, with only 128 and 192 tokens per image instead of the original 576, our approach nearly retains the performance of the original LLaVA-1.5, with only negligible performance declines of 1.0% and 1.2%, respectively. Our approach becomes even more advantageous with higher reduction rates: With an aggressive 88.9% reduction rate, our approach results in only a 3.3% degradation in average performance across benchmarks, outperforming the second-best VisionZip Yang et al. [2025] by a substantial margin of 4.0%. These results suggest that our method effectively maintains high performance while reducing the number of visual tokens processed.
@@
Results on LLaVA-NeXT. In Table 3, we further compare the performance achieved by our approach with other state-of-the-arts on the more advanced LLaVA-NeXT-7B Liu et al. [2024b]. We present the performance of all approaches under a fixed budget of 320 tokens per image, corresponding to an 88.9% reduction rate. Our approach continues to achieve superior performance on LLaVA-NeXT-7B Liu et al. [2024b], attaining the best performance on 6 out of 9 benchmarks, and achieving 95.4% of the original LLaVA-NeXT-7B Liu et al. [2024b] performance with only 11.1% of the token budget without additional training.
@@
Results on Qwen-2.5-VL. To further validate the general effectiveness of our approach, we apply it to the more advanced Qwen-2.5-VL Bai et al. [2025] with three different LLM scales and visualize the performance on three general visual understanding benchmarks in Figure 5. The full numerical results are also provided in Appendix A.7. As shown, our approach consistently outperforms FastV Chen et al. [2024a], PDrop Xing et al. [2025] and VisionZip Yang et al. [2025] across all retention rates (from 11.1% to 50%) and model scales. We also observe that on MME, both FastV and PDrop exhibit degraded performance under low token budgets as the LLM scales from 7B to 32B. We speculate that this is due to larger LLMs introducing stronger language priors, which can hinder the accurate selection of critical visual tokens. In contrast, our two-stage reduction strategy mitigates this bias and remains robust across LLM scales, demonstrating superior performance.
@@
Table 4: Performance comparisons on Qwen-2.5-VL-7B Bai et al. [2025] across 3 referring expression comprehension benchmarks: RefCOCO, RefCOCO++, and RefCOCOg. The best results in each setting are bolded, and the second-best are underlined. Evaluation is based on our re-implementation.
Method Venue RefCOCO RefCOCO++ RefCOCOg Average
@@
[…]
Qwen-2.5-VL-7B Bai et al. [2025] arXiv’25 89.45 92.56 85.16 83.50 89.02 79.15 86.76 87.24 100.0%
@@
FastV Chen et al. [2024a] ECCV’24 85.27 87.84 82.28 79.02 82.95 72.86 82.95 83.32 94.8%
PyramidDrop Xing et al. [2025] CVPR’25 87.79 91.00 83.22 81.48 86.55 74.02 84.62 85.10 97.2%
VisionZip Yang et al. [2025] CVPR’25 88.33 91.55 83.68 81.99 87.28 74.09 85.08 85.92 97.9%
@@
FastV Chen et al. [2024a] ECCV’24 73.85 73.38 74.21 66.75 68.88 62.65 71.06 71.86 81.2%
PyramidDrop Xing et al. [2025] CVPR’25 77.52 80.82 72.07 70.27 75.48 63.33 74.86 75.65 85.1%
VisionZip Yang et al. [2025] CVPR’25 80.92 84.55 76.91 74.02 79.52 67.89 78.12 79.34 89.7%
@@
FastV Chen et al. [2024a] ECCV’24 43.57 46.81 40.86 39.47 43.78 36.02 43.04 42.69 48.5%
PyramidDrop Xing et al. [2025] CVPR’25 46.46 53.83 37.23 42.29 47.76 32.81 45.32 44.91 50.4%
VisionZip Yang et al. [2025] CVPR’25 60.12 66.95 55.41 58.33 64.88 48.52 57.09 56.44 67.8%
@@
We extend our comparisons to more challenging grounding tasks and present the performance of different approaches in Table 4. Compared to image understanding tasks, these grounding tasks require higher token budgets to preserve visual information necessary for precise localization. A 75% reduction rate, for instance, halves the performance of FastV Chen et al. [2024a] and PDrop Xing et al. [2025]. In this challenging scenario, our approach still robustly maintains 80.7% of the original performance. Additionally, our approach achieves 96.1% of the original performance with only 50% of the visual tokens. Qualitative results in Figure A2 further demonstrate that our approach effectively retains the critical tokens for accurate localization in challenging cases.
Method TGIF MSVD MSRVTT ActivityNet
Acc. Score Acc. Score Acc. Score Acc. Score
Video-LLaVA-7B Lin et al. [2024] 47.0 3.40 70.5 3.92 58.3 3.51 42.2 3.37
FastV Chen et al. [2024a] 42.7 3.19 67.4 3.83 53.6 3.40 36.1 3.15
PyramidDrop Xing et al. [2025] 44.1 3.26 66.7 3.81 56.1 3.45 37.4 3.15
VisionZip Yang et al. [2025] 44.0 3.28 68.2 3.85 55.4 3.42 39.0 3.23
VScan (Ours) 46.9 3.35 69.8 3.93 57.1 3.48 42.6 3.34
Table 5: Performance comparisons on Video-LLaVA-7B Lin et al. [2024] across 4 video understanding tasks with a 75% reduction rate. The best results are bolded, and the second-best are underlined. Evaluation is based on our re-implementation.
Results on Video-LLaVA. Finally, we validate the effectiveness of our approach on video understanding tasks and compare its performance against other approaches on Video-LLaVA-7B Lin et al. [2024], as shown in Table 5. Specifically, we report the accuracy and GPT-evaluated scores for each benchmark to assess the quality of the responses. The experimental results demonstrate that, even under a strict 25% token budget, our approach is able to preserve nearly the full performance of the original Video-LLaVA-7B. Moreover, it consistently surpasses prior methods across multiple benchmarks, highlighting the efficiency and robustness of our token reduction strategy in challenging video understanding tasks.
Additional Results. In Appendix A, we present further experimental results, demonstrating that our VScan generalizes to more challenging benchmarks that require fine-grained visual perception (e.g., DocVQA Mathew et al. [2021] and InfoVQA Mathew et al. [2022]) and also effectively accelerates training while largely preserving model accuracy.
@@ 5.3 Ablation Studies
Varying Retention Rates $R_{1}$ and $R_{2}$. We analyze how different retention rate configurations affect performance by varying the retention rates $R_{1}$ and $R_{2}$ during the visual encoding and LLM decoding stages, respectively. As shown in Table 6 (a), we observe that relying solely on token selection in the visual encoder or applying overly aggressive pruning during LLM decoding leads to suboptimal performance. Specifically, text-agnostic pruning ($R_{1}$) reduces computation early but lacks awareness of the text prompt, making it prone to removing visually subtle yet instruction-critical cues. In contrast, text-aware pruning ($R_{2}$) benefits from intermediate cross-modal alignment and can better retain tokens relevant to the query, though pruning too aggressively at this stage may interrupt the model’s semantic refinement and harm reasoning stability. Instead, a more balanced and gradual two-stage pruning strategy, i.e., $R_{1}=16.7\%$ and $R_{2}=33.3\%$, achieves the best performance. These results also validate the effectiveness of combining both visual token reduction strategies to jointly improve efficiency and maintain accuracy.
@@
Different Encoding Layers for Local Scan. In Table 6 (c), we explore the effect of performing local token selection at different layers of the visual encoder. Consistent with our empirical findings in Section 3, we find that applying the local scan at a shallow layer ($l=6$) yields the best performance. However, performing the local scan at very early ($l=2$) or the output layer ($l=23$) leads to a noticeable performance drop, with performance on GQA falling to 57.1 and 57.3, respectively. This confirms that local tokens selected from shallow layers are more effective at capturing fine-grained visual details and better complement the global token set.